# SpecEval: Evaluating Code Comprehension in Large Language Models via Program Specifications

### v1

@@ 4.1. Benchmark Construction
We construct a benchmark that includes programs and corresponding ground-truth specifications for evaluation. Concerning the specification language, we adopt the Java Modeling Language (JML) (Leavens et al., 1998) for Java programs for evaluation since they can be conveniently validated by OpenJML (Cok, 2011). We use the released datasets by the latest two research of LLM-based specification generation (Wen et al., 2024; Ma et al., 2024) for evaluation. One work by Ma et al. (Ma et al., 2024) utilizes two datasets, SpecGenBench and the Java Category of SV-COMP Benchmark (sosy lab, 2024). The former contains 120 Java programs with manually-written strong specifications and the latter is widely adopted in the evaluation of software verification tools. For SpecGenBench, we extract 117 programs from the dataset along with corresponding specifications, whereas 3 programs are discarded due to incompatibility with OpenJML. For the Java Category of SV-COMP Benchmark, as these programs are without corresponding specifications, we manually craft them following the similar procedure adopted in the construction of SpecGenBench (Ma et al., 2024). Extensive efforts have been conducted to manually write these specifications, and we have succeeded in 47 programs from SV-COMP that are compatible with OpenJML, and the written specifications can fully specify the program semantics. Another work (Wen et al., 2024) leverages another dataset, Frama-C-Problems (github, 2024), which contains 47 C programs and corresponding ACSL (Baudin et al., 2008) specifications. To cater to Java style, we manually rewrote the programs in Java and altered the specifications to JML-style. 40 programs that can be equivalently translated into Java programs. Others are discarded due to specific features (i.e., pointers) of the C programming language that cannot be trivially translated.
@@
All specifications have been cross-validated by three experts in formal methods and programming, where the semantics of the specifications and corresponding programs are compared. The specifications will be re-formulated until their strength is confirmed by all the experts. This is to ensure high quality within the specifications and ability to describe the behaviors and functionalities of the target program thoroughly. Additionally, all specifications are also ensured to be verifiable, i.e., able to pass the verification of the verifier (namely OpenJML (Cok, 2011) in our work), so the correctness of the specifications can be guaranteed. Eventually, 204 programs are collected for evaluation.
@@
During the experiment, we adopt dynamic validation supported by specification verifiers to check the correctness of the specifications generated by subject LLMs (i.e., to calculate $\mathbf{Cr}(s)$ mentioned in Section 3.2 for a given specification $s$). A specification is considered correct (i.e., $\mathbf{Cr}(s)=\mathbf{true}$) if it maintains its consistency with the target program during a runtime execution process, where test cases are utilized to invoke the target programs. To this end, test cases are also prepared for each program. Similar to the preparation of ground-truth specifications, we inherit existing test cases in the original programs if there are or employ experts to manually craft them otherwise. The experts are required to cover as many branches as possible with the written test cases. False specifications (those in Section 3.2.1 and Section 3.2.2) are generated to test whether the test cases can disprove them, and more test cases should be crafted if they failed. Each program receives, on average, 15.10 test cases, with an average branch coverage of 95.54% and line coverage of 93.34%. The average LoC (Lines of Code) and CC (Cyclomatic Complexity) of programs in the benchmark are 20.06 and 6.34, respectively. Detailed statistics of the benchmark are provided on our website (Author, s).
@@ 4.2. Studied LLMs
In this work, we selected six state-of-the-art and popular LLMs for evaluation.
  • GPT-3.5-Turbo (OpenAI, 2023), a general LLM developed by OpenAI (OpenAI, 2024a), able to understand and generate natural language or code. We adopt gpt-3.5-turbo-0125, the latest model in the GPT-3.5 family during the experiment, for evaluation.
  • GPT-4-Turbo (Achiam et al., 2023), the latest general LLM developed by OpenAI, with the most state-of-the-art performance among all general LLMs currently.
  • Llama 2 (Touvron et al., 2023), an open-source general LLM developed by Meta AI (AI, 2024). We adopt Llama-2-7b-chat for evaluation.
  • CodeLlama (Roziere et al., 2023), an open-source code LLM based on the architecture of Llama 2. We adopt CodeLlama-7b-Instruct for evaluation.
@@
  • Magicoder (Wei et al., 2023), an open-source code LLM fine-tuned with OSS-Instruct, a novel high-quality code instruction data. The version employed in the evaluation is Magicoder-S-DS-6.7B.
The selected models have been applied to a wide variety of code-related tasks in existing works (Chen et al., 2021; Liu et al., 2023; Deng et al., 2023). They also exhibit diversities over multiple dimensions, ranging from general LLMs to code LLMs, open-source models to closed-source models, and foundation models to fine-tuned models. Due to the computing resource limit, only 7B versions of open-source LLMs are selected.
@@ 4.3. Prompt Format and Model Configurations
In terms of the prompts provided to LLMs, we generally adopt a unified prompt structure for each task and all the subject LLMs in the experiments. The structure mainly consists of four components. Initially, a system message is appended to the context to inform the model of its basic role. The few-shot prompting technique is then adopted, where several pairs of example requests and replies are provided to the LLM, assisting the LLM in learning the task requirements and desired output format. Afterward, a description of the target task (i.e., the current task that the LLM is expected to answer) is appended to the context. Up to this point, the context constitutes a complete query and is sent to the LLM, which learns necessary information from the context and completes the task within its reply. The template prompts for different tasks are provided on our website (Author, s).
For the open-source models mentioned in Section 4.2, we deploy a local server to execute the models and process relevant queries, running on a Linux Server equipped with one H800 GPU with 80GB memory. The source code and model weights of all open-source models can be accessed at corresponding HuggingFace (HuggingFace, 2024) repositories. For the closed-source models (i.e., GPT-3.5 and GPT-4), we access them through the API provided by OpenAI (OpenAI, 2024b). Following the settings of previous works (Du et al., 2024; Liu et al., 2024a), we adopt greedy decoding (Chen et al., 2015) for all subject LLMs to make the results reproducible. Generally, the models are configured as top_k=1 and temperature=0 under greedy decoding. The max context length is set to 2048 tokens, leaving abundant space for the LLMs to generate detailed results. Other parameters are kept consistent with default model settings.
@@ 5. Experimental Results
We aim to answer the following research questions through evaluation:
  • RQ1: How is the performance of LLMs on specification-related tasks generally? This gives us an overall scenario for the performance of LLMs on each task.
  • RQ2: How is the performance of LLMs against semantic-preserving perturbations? This evaluates the impact of the perturbations through Counterfactual Analysis.
  • RQ3: Can LLMs maintain performance consistency over progressive tasks? It assesses the consistency of LLMs over different tasks through Progressive Consistency Analysis.
@@ 5.1.1. Comparison Between Tasks
For different tasks involved in the experiments, the models showcase different levels of fulfillment. We take the original category as the example for illustration. For the Candidate Judgement task, the majority of models can achieve an accuracy above 50%, with some excelling models achieving relatively high accuracy, such as the 81.37% of GPT-4. For the Candidate Selection task, the accuracy of most models shows some decline compared to that of Judgement (from 57.84% to 29.41% of CodeLlama) or has no significant variance (from 44.12% to 39.71% of Magicoder). When it comes to the Specification Infilling task, we can witness a significant drop in the performance of all models among all categories, ranging from 7.84% of Llama to 31.86% of Magicoder. As for the Specification Generation task, the average precision of most models ranges from 24.01% of Llama to 54.05% of GPT-3.5, whereas the average recall is relatively lower, varying between 15.88% of Llama and 37.80% of GPT-3.5. The low recall of all LLMs suggests that they can only generate limited specifications with strength comparable to the ground truth, indicating a relatively weak ability to summarize full program semantics into the form of formal specifications. The tasks for the other five categories showcase similar performance variance patterns as well.
@@
According to their performance, the most straightforward task involved in the experiment is Judgement, followed by Selection, then Generation, whereas the most challenging task is Infilling. It is an interesting finding, and to our intuition, the root cause of this phenomenon lies in the different aspects of ability that are examined by different tasks. Judgement and Selection are mainly about reading and understanding (i.e., ”input”) specifications, where LLMs try to comprehend given specifications and give relevant conclusions. On the contrary, Generation is about writing and expressing (i.e., ”output”) specifications, where LLMs can freely articulate code semantics they understand in the form of specifications. The criteria for successful generation is to pass the verifier. Thus, the written specifications may not be strong. The tasks mentioned above examine two opposite aspects of ability concerning specifications. Interestingly, Infilling lies somewhere in between, where LLMs write partial specifications but cannot do it at any will. They have to generate results within the structure defined by the target specification to fill in, which are usually strong with sufficient semantic expression. To this end, they have to read and understand the given specification first; some may be difficult for them, and then, based on their understanding, write the standard answers. Consequently, the Infilling task examines both of the two aspects mentioned above (i.e., ”input” and ”output”) simultaneously, making it the most difficult task of all.
Acc. Acc. Acc. #Pass AvgPrec. AvgRec. #Pass Acc. Acc. Acc. #Pass AvgPrec. AvgRec. #Pass
Original GPT-3.5 64.71 67.65 23.53 48 54.05 37.80 79 Def-use Break GPT-3.5 67.65 61.27 27.45 56 53.37 38.47 77
GPT-4 81.37 89.71 11.27 23 32.90 22.75 54 GPT-4 78.92 86.27 14.22 29 24.28 16.91 34
Llama 50.00 30.88 7.84 16 24.01 15.88 21 Llama 54.41 24.02 5.39 11 20.98 16.02 16
CodeLlama 57.84 29.41 17.65 36 43.70 26.05 61 CodeLlama 62.25 28.92 15.69 32 35.32 23.04 43
Deepseek-coder 57.35 61.27 17.65 36 43.38 31.37 55 Deepseek-coder 60.29 59.80 10.29 21 40.79 29.99 46
Magicoder 44.12 39.71 31.86 65 53.22 34.95 67 Magicoder 53.43 45.59 31.86 65 48.44 35.44 54
If-else Flip GPT-3.5 68.14 65.20 22.55 46 42.61 31.26 62 Independent Swap GPT-3.5 66.18 66.18 25.00 51 51.53 39.02 75
GPT-4 77.94 83.82 11.76 24 20.20 14.20 28 GPT-4 79.90 86.27 12.75 26 23.78 18.23 32
Llama 49.51 24.51 8.82 18 28.59 21.71 23 Llama 48.53 22.55 7.84 16 24.53 17.82 20
CodeLlama 63.24 30.88 13.24 27 38.28 27.79 43 CodeLlama 65.69 35.29 14.71 30 36.26 25.38 42
Deepseek-coder 61.27 57.35 17.16 35 41.79 30.93 55 Deepseek-coder 55.88 59.80 18.14 37 42.07 31.03 55
Magicoder 55.88 50.00 32.35 66 47.33 35.95 53 Magicoder 58.82 43.14 35.29 72 44.35 32.85 52
Name Random GPT-3.5 63.24 54.90 24.02 49 48.49 37.16 74 Name Shuffle GPT-3.5 68.14 66.67 20.59 42 41.43 30.99 57
GPT-4 76.96 81.37 10.78 22 25.55 18.57 40 GPT-4 76.96 84.31 12.75 26 29.52 21.26 46
Llama 56.37 25.00 10.78 22 22.79 16.45 20 Llama 49.02 26.96 7.35 15 28.88 19.77 31
CodeLlama 59.31 25.00 15.20 31 38.35 25.00 53 CodeLlama 59.80 30.39 15.69 32 33.47 24.48 37
Deepseek-coder 55.88 61.27 18.63 38 37.54 28.32 45 Deepseek-coder 59.80 59.31 13.73 28 36.75 28.63 41
Magicoder 60.29 48.04 36.76 75 43.79 34.73 54 Magicoder 63.73 46.08 28.43 58 45.57 35.03 51
@@ 5.1.2. Comparison Between Models
It can be observed that different models are specialized in different tasks. Regarding Judgement, GPT-4 outperforms all other models significantly with leading accuracy across all categories, followed by GPT-3.5, of which the accuracy is ranked second in all categories. The three code LLMs involved in the experiments underperform the GPT family on this task. Regarding Selection, a similar pattern can be observed across all categories, where GPT-4 is way ahead, followed by GPT-3.5, then the three code LLMs. It is also worth noting that Llama and CodeLlama, two models with the same architecture, suffer from significant performance declines in Selection tasks compared to Judgement among all categories. As described in Section 5.1.1, Judgement and Selection are mainly about reading and understanding specifications. This indicates that GPT-4 and GPT-3.5 showcase impressive ability regarding the comprehension of specifications.
Nevertheless, the ranking varies significantly when it comes to the latter two tasks. In terms of Infilling, the best-performing model is Magicoder, the only model achieving an accuracy above 30% across 5 out of 6 categories. The following model is GPT-3.5, another model that can keep an accuracy above 20% among all categories. Other models demonstrate a relatively lower performance, with accuracy below 20% in all categories. In terms of Generation, the leading model is GPT-3.5, with the highest average precision and recall over 4 out of 6 categories, along with the highest #Pass over all 6 categories. The second-best model is Magicoder, able to surpass GPT-3.5 in 2 out of 6 categories concerning precision and recall.
Thus, we can find that, despite the excelling performance of the closed-source GPT models by OpenAI, it is still possible for open-source models like Magicoder to achieve comparable or even better performance with a much smaller model scale under the assistance of fine-tuning with high-quality code data corpora.
Answer to RQ1: LLMs showcase a limited ability to articulate program semantics with formal specifications. The best-performing task is Judgement, followed by Selection, then Generation, and the most challenging task is Infilling. Notably, Magicoder exhibit leading performance in Infilling and Generation, indicating the possibility for small-scale LLMs to surpass large-scale ones.

### v2

@@ 4.1. Benchmark Construction
We construct a benchmark that includes programs and corresponding ground-truth specifications for evaluation. Concerning the specification language, we adopt the Java Modeling Language (JML) (Leavens et al., 1998) for Java programs for evaluation since they can be conveniently validated by OpenJML (Cok, 2011). We use the released datasets by the latest two works on LLM-based specification generation (Wen et al., 2024; Ma et al., 2024). One work by Ma et al. (Ma et al., 2024) utilizes two datasets, SpecGenBench and the Java Category of SV-COMP Benchmark (sosy lab, 2024). The former contains 120 Java programs with manually-written strong specifications and the latter is widely adopted in the evaluation of software verification tools. For SpecGenBench, we extract 117 programs from the dataset along with corresponding specifications, whereas 3 programs are discarded due to incompatibility with OpenJML. For the Java Category of SV-COMP Benchmark, as these programs are without corresponding specifications, we manually craft them following the similar procedure adopted in the construction of SpecGenBench (Ma et al., 2024). Extensive efforts have been conducted to manually write these specifications, and we have succeeded in 47 programs from SV-COMP that are compatible with OpenJML, and the written specifications can fully specify the program semantics. Another work (Wen et al., 2024) leverages another dataset, Frama-C-Problems (github, 2024), which contains 47 C programs and corresponding ACSL (Baudin et al., 2008) specifications. To cater to Java style, we manually rewrote the programs in Java and altered the specifications to JML-style. 40 programs that can be equivalently translated into Java programs. Others are discarded due to specific features (i.e., pointers) of the C programming language that cannot be trivially translated.
@@
All specifications have been cross-validated by three experts in formal methods and programming, where the semantics of the specifications and corresponding programs are compared. The specifications will be re-formulated until their strength is confirmed by all the experts to ensure their quality and ability to describe the behaviors and functionalities of the target program thoroughly. Additionally, all specifications are also ensured to be verifiable, i.e., able to pass the verification of the verifier (namely OpenJML (Cok, 2011) in our work), so the correctness of the specifications can be guaranteed. Eventually, 204 programs are collected for evaluation. The average LoC (Lines of Code) and CC (Cyclomatic Complexity) of the programs adopted are 20.06 and 6.34, respectively. Among the programs, 36 are sequential (i.e., no branches or loops), 77 are loop-free but with branches (i.e., if-else or switch statements), 70 have single layer loops, and 20 have nested loops.
@@
During the experiment, we adopt runtime checking supported by specification verifiers to validate the correctness of the specifications generated by the studied LLMs (i.e., to calculate $\mathbf{Cr}(s)$ mentioned in Section 3.1 for a given specification $s$). A specification is considered correct (i.e., $\mathbf{Cr}(s)=\mathbf{true}$) if it maintains its consistency with the target program during a runtime execution process. To this end, test cases are also prepared for each program to invoke their execution. Similar to the preparation of ground-truth specifications, we inherit existing test cases in the original programs if there are or employ experts to manually craft them otherwise. The experts are required to cover as many branches as possible with the written test cases. False specifications (those in Section 3.1.1 and Section 3.1.2) are generated to test whether the test cases can disprove them, and more test cases should be crafted if they failed. Each program receives, on average, 15.10 test cases, with an average branch coverage of 95.54% and line coverage of 93.34%.
@@ 4.2. Studied LLMs
In this work, we selected six state-of-the-art LLMs for evaluation.
  • GPT-3.5-Turbo (OpenAI, 2023), a general LLM developed by OpenAI (OpenAI, 2025a). We adopt gpt-3.5-turbo-0125, the latest model in the GPT-3.5 family during the experiment, for evaluation.
  • GPT-4-Turbo (Achiam et al., 2023), an improved general LLM developed by OpenAI, with leading performance among all general LLMs currently.
  • Llama 2 (Touvron et al., 2023), an open-source general LLM developed by Meta AI (AI, 2025). We adopt Llama-2-7b-chat for evaluation.
  • CodeLlama (Roziere et al., 2023), an open-source code LLM based on the architecture of Llama 2. CodeLlama-7b-Instruct is adopted.
@@
  • Magicoder (Wei et al., 2023), an open-source code LLM fine-tuned with OSS-Instruct, a novel high-quality instruction tuning technique. The version employed in the evaluation is Magicoder-S-DS-6.7B.
The selected models have been applied to a wide variety of code-related tasks in existing works (Chen et al., 2021; Liu et al., 2023a; Deng et al., 2023). They also exhibit diversities over multiple dimensions, ranging from general LLMs to code LLMs, open-source models to closed-source models, and foundation models to fine-tuned models. Due to the computing resource limit, only 7B versions of open-source LLMs are selected.
@@ 4.3. Prompt Format and Model Configurations
We generally adopt a unified prompt structure for each task and all the subject LLMs in the experiments. The structure mainly consists of four components. Initially, a system message is appended to the context to inform the model of its basic role. The few-shot prompting technique is then adopted, where several pairs of example requests and replies are provided to the LLM, assisting the LLM in learning the task requirements and desired output format. In our experiments, a 2-shot setting is adopted. The few-shot examples are collected from the official tutorial of JML (OpenJML, 2025), including example specifications that cover the fundamental syntax and semantics of JML, enough for LLMs to understand and accomplish the tasks. Afterward, a description of the target task (i.e., the current task that the LLM is expected to answer) is appended to the context. Up to this point, the context constitutes a complete query and is sent to the LLM, which learns necessary information from the context and completes the task within its reply.
For the open-source models mentioned in Section 4.2, we deploy a local server to execute the models and process relevant queries, running on a Linux Server equipped with one H800 GPU with 80GB memory. The source code and model weights of all open-source models can be accessed at corresponding HuggingFace (HuggingFace, 2025) repositories. For GPT-3.5 and GPT-4, we access them through the API provided by OpenAI (OpenAI, 2025b). Following the settings of previous works (Du et al., 2024; Liu et al., 2024a), we adopt greedy decoding (Chen et al., 2015) for all subject LLMs to make the results reproducible. Generally, the models are configured as top_k=1 and temperature=0 under greedy decoding. Other parameters are kept consistent with default settings.
Acc. Acc. Acc. #Pass Prec. Rec. #Pass Acc. Acc. Acc. #Pass Prec. Rec. #Pass
Original GPT-3.5 64.71 67.65 34.31 70 73.22 25.35 106 Def-use Break GPT-3.5 67.65 61.27 36.76 75 70.51 29.82 98
GPT-4 81.37 89.71 41.18 84 83.37 34.17 135 GPT-4 78.92 86.27 40.20 82 77.81 46.03 112
Llama 50.00 30.88 10.78 22 29.65 4.07 24 Llama 54.41 24.02 6.86 14 24.68 3.26 20
CodeLlama 57.84 31.37 21.08 43 55.22 10.77 74 CodeLlama 62.25 28.43 19.61 40 46.01 10.77 55
Deepseek-coder 57.35 61.27 29.90 61 61.56 17.39 75 Deepseek-coder 60.29 59.80 24.51 50 61.80 23.38 72
Magicoder 44.12 39.71 29.41 60 64.16 13.72 81 Magicoder 53.43 45.59 29.41 60 55.53 19.76 60
If-else Flip GPT-3.5 68.14 65.20 36.27 74 66.99 34.61 97 Independent Swap GPT-3.5 66.18 66.18 37.75 77 75.32 38.01 112
GPT-4 77.94 83.82 40.69 83 79.16 48.26 115 GPT-4 79.90 86.27 41.67 85 77.69 47.55 115
Llama 49.51 24.51 12.25 25 33.22 7.31 25 Llama 48.53 22.55 11.27 23 31.63 5.74 26
CodeLlama 62.25 31.37 17.65 36 49.00 16.36 56 CodeLlama 64.22 31.86 20.59 42 48.39 14.84 54
Deepseek-coder 61.27 57.35 27.45 56 63.26 27.27 79 Deepseek-coder 55.88 59.80 30.88 63 63.43 25.62 78
Magicoder 55.88 50.00 30.39 62 55.88 18.61 62 Magicoder 58.82 43.14 34.80 71 57.50 23.18 68
Name Random GPT-3.5 63.24 54.90 35.78 73 71.65 38.14 103 Name Shuffle GPT-3.5 68.14 66.67 33.82 69 68.27 29.91 97
GPT-4 76.96 81.37 35.29 72 81.18 50.65 124 GPT-4 76.96 84.31 35.78 73 77.39 45.69 111
Llama 56.37 25.49 10.29 21 28.73 5.70 25 Llama 49.02 26.96 8.82 18 32.64 5.09 31
CodeLlama 61.27 24.51 19.61 40 49.10 12.75 65 CodeLlama 61.27 30.88 22.06 45 43.70 13.06 48
Deepseek-coder 55.88 61.27 27.45 56 61.71 26.86 69 Deepseek-coder 59.80 59.31 25.49 52 61.40 27.58 70
Magicoder 60.29 48.04 37.25 76 52.04 22.24 63 Magicoder 63.73 46.08 28.92 59 54.17 21.37 63
@@ 5. Experimental Results
Our experiments aim to answer the following research questions:
  • RQ1: How is the general performance of each LLM evaluated on the tasks related to formal specifications?
  • RQ2: How does few-shot prompting strategy affect the performance of LLMs in the tasks involved?
  • RQ3: How is the performance variance of LLMs when confronted with semantic-preserving perturbations?
@@ 5.1.1. Comparison Between Tasks
For different tasks involved in the experiments, the models showcase different levels of fulfillment. We take the original category as the example for illustration. For the Judgement task, most models can achieve an accuracy below 65% except the 81.37% of GPT-4. Given the high similarity between this task and the model checking technique, the models may lack the symbolic reasoning capabilities required for model checking. For the Selection task, the accuracy of most models either shows some decline compared to that of Judgement or has no significant variance. When it comes to Infilling, we can witness a significant drop in the performance of all models. As for Generation, the average precision ranges from 29.65% of Llama to 83.37% of GPT-4, whereas the average recall is much lower, varying between 4.07% of Llama and 34.17% of GPT-4. The low recall of all LLMs suggests that they can only generate limited specifications with strength comparable to the ground truth, indicating a relatively weak ability to summarize full program semantics into formal specifications. The other five categories also display similar patterns.
@@
According to their performance, the most straightforward task involved in the experiment is Judgement, followed by Selection, then Generation, whereas the most challenging task is Infilling. It is an interesting finding, and to our intuition, the root cause of this phenomenon lies in the different aspects of ability that are examined by different tasks. Judgement and Selection are mainly about reading and understanding (i.e., ”input”) specifications, where LLMs try to comprehend given specifications and give relevant conclusions. On the contrary, Generation is about writing and expressing (i.e., ”output”) specifications, where LLMs can freely articulate code semantics they understand in the form of specifications, and the specifications written may not be strong. Interestingly, Infilling lies somewhere in between, where LLMs write partial specifications but cannot do it at any will. They have to generate results within the structure defined by the target specification to fill in, which is usually strong with sufficient semantic information. To this end, they have to read and understand the given specification first, which may be difficult for them, and then, based on their understanding, figure out the desired answers. Consequently, the Infilling task examines both of the two aspects mentioned above (i.e., ”input” and ”output”) simultaneously, making it the most difficult task of all.
Table 2. Performance of all models under different few-shot settings on the Original category. Config: The adopted few-shot configuration. Acc.: Average Accuracy, Prec.: Average Precision, Rec.: Average Recall (all in percentages).
Model Config Judgment Selection Infilling Generation Model Config Judgment Selection Infilling Generation Model Config Judgment Selection Infilling Generation
Acc. Acc. Acc. Prec. Rec. Acc. Acc. Acc. Prec. Rec. Acc. Acc. Acc. Prec. Rec.
GPT-3.5 0-shot 62.25 25.49 29.90 34.83 7.30 GPT-4 0-shot 77.94 25.00 36.76 83.64 32.73 Llama 0-shot 42.65 25.00 20.10 0.98 0.00
1-shot 64.71 62.75 26.96 59.66 15.87 1-shot 81.37 85.29 37.75 79.56 31.23 1-shot 56.86 26.96 14.22 25.63 2.00
2-shot 64.71 67.65 34.31 73.22 25.35 2-shot 81.37 89.71 41.18 83.37 34.17 2-shot 50.00 30.88 14.22 24.01 15.88
CodeLlama 0-shot 52.45 23.53 3.43 6.21 1.82 Deepseek Coder 0-shot 53.43 25.00 14.22 63.58 6.76 Magicoder 0-shot 55.39 23.04 14.71 69.90 5.71
1-shot 54.22 27.94 14.12 38.61 10.37 1-shot 55.88 51.96 28.92 61.91 14.40 1-shot 62.75 42.16 18.14 48.40 9.48
2-shot 57.84 31.37 21.08 55.22 10.77 2-shot 57.35 61.27 29.90 61.56 17.39 2-shot 44.12 39.71 29.41 64.16 13.72
@@ 5.1.2. Comparison Between Models
Different models are specialized in different tasks. Regarding Judgement and Selection, GPT-4 outperforms all other models significantly with leading accuracy across all categories, followed by GPT-3.5. The three code LLMs involved in the experiments yield a relatively low accuracy. As described in Section 5.1.1, Judgement and Selection are mainly about reading and understanding specifications. This indicates that GPT-4 and GPT-3.5 showcase impressive ability regarding the comprehension of specifications. In terms of Infilling, the gap between GPT-4 and GPT-3.5 is significantly narrower as task difficulty rises. Notably, Magicoder demonstrates the most outstanding performance among the open-source models, which even surpasses the GPTs in the Name Random category. This indicates the possibility for open-source models to achieve comparable or even better performance compared to the closed-source GPT models with a much smaller model scale, under the assistance of fine-tuning with high-quality code data corpora. In the Generation task, the models’ ranking remains largely consistent, except for DeepSeek-Coder emerging as the top-performing model among the open-source ones.
Figure 3. The selection task for program FizzBuzz, the correct answer, and the options chosen by each model.
[…]
Answer to RQ1: LLMs showcase a limited ability to specify program semantics with formal specifications. Furthermore, LLMs are relatively weak at synthesizing together the semantics of multiple control-flow structures. Extraneous information with unknown correctness in prompts also poses a notable negative impact on the models’ reasoning abilities, especially when correct and incorrect information are syntactically and semantically similar.
5.2. RQ2: Effectiveness of the Few-shot Setting
As is described in Section 4.3, we adopted few-shot prompts for all models. To investigate the effect of few-shot prompts on the models’ comprehension abilities, and validate the effectiveness of our experimental design, we conduct further experiments by setting up 0-shot and 1-shot configurations in comparison with the 2-shot setting adopted in Section 5.1.
Table 2 shows the performance of all models under different few-shot settings. Generally, the effect of few-shot learning is obvious, with different degrees of improvement on most tasks across all models. For Judgement, the improvement of few-shot prompting is relatively modest, bringing less than 10% improvement for all models. For Selection, dramatic improvement can be observed on GPT-3.5, GPT-4, and Deepseek-Coder. As for Infilling, the 2-shot setting can increase the accuracy of open-source models from around 10% to around 30% compared to 0-shot. For Generation, the 2-shot configuration significantly improves the precision of GPT-3.5, Llama, and CodeLlama, while also enhancing the recall of all models to some extent. We did not conduct further experiments on configurations providing more few-shot examples due to frequent context length overflows and GPU memory shortages (with only one H800 GPU for experiments) caused by such configurations. Since we aim to evaluate model performance under a consistent setting to avoid bias rather than study the best setting for each model, the 2-shot setting is effective and fair. Some models exhibit slight performance regression, such as the precision of GPT-4 in Generation. This is due to the model’s tendency to generate additional content to facilitate reasoning when no examples are provided to rectify its output. Since some models may not have been previously trained on datasets containing extensive specifications, few-shot learning can provide them with the basic syntax and semantics of specification languages (i.e., JML in our experiments) within the context, equipping them with the fundamental knowledge needed to accomplish the experimental tasks. For these models, few-shot learning can significantly enhance their performance. For models already equipped with relevant knowledge, the assistance from few-shot learning may be relatively limited, but the provided few-shot examples can still achieve some improvements by familiarizing the models with the task content and rectifying their output format.
Answer to RQ2: Few-shot prompting can effectively improve the performance of all models across all tasks, especially when the models are not equipped with the basic knowledge of the specification language involved, where few-shot examples effectively help models learn specification syntax and semantics, validating the effectiveness of the 2-shot configurations in our experimental settings.